How many moles of precipitate are formed when 24.0 mL of 0.250 M Ca(NO₃)₂ is mixed with excess Li₃PO₄ in the following chemical reaction?
3 Ca(NO₃)₂ (aq) + 2 Li₃PO₄ (aq) → Ca₃(PO₄)₂ (s) + 6 LiNO₃ (aq)
1
Expert's answer
2020-03-30T10:50:41-0400

"3 Ca(NO\u2083)\u2082 (aq) + 2 Li\u2083PO\u2084 (aq) \u2192" "Ca\u2083(PO\u2084)\u2082 (s) + 6 LiNO\u2083 (aq)"

As "Li_3PO_4" is in excess,reaction is independent of its concentration.

As per the reaction,"3" moles of calcium nitrate produce "1" mole of calcium phosphate or

The precipitate formed will be one-third of the reactant.

Moles of "Ca(NO_3)_2=M\times V (in\ L)="

"0.25\times24\times 10^{-3}=0.006\ mol"

Moles of "Ca_3(PO_4)_2" formed "=\frac{1}{3}\times 0.006=0.002\ mol"